# Application of Operator Theory for the Collatz Conjecture

### v1

@@ Abstract
In this paper, we formulate the Collatz conjecture (or the $13n{+}$ 1-problem) as some operator theoretic problems and prove that each of those statements is equivalent to the conjecture.
@@ 1 Introduction
Throughout this paper, $\mathbb{N}$ denotes the set of all positive integers and $\left\lvert x\right\rvert$ denotes the absolute value of a number $x$. For a topological space $X$ and its subspace $Y$, $\overline{Y}$ denotes the closure of $Y$ in $X$.
@@ 2.1 Dynamical systems
$P(x)=f^{\tau(x)}(x),\text{ where }x\in\Sigma^{\prime}.$ (2)
@@ 2.2 The Collatz conjecture
$f(n)=\begin{cases}3n+1,&n:\text{odd},\\ n/2,&n:\text{even}.\end{cases}$ (3)
@@ 2.3 Equivalence relations
We define a binary relation on $X$ which is called the orbit equivalence relation by $f$ as follows:
$x\sim y\iff\operatorname{orb}(x;f)\cap\operatorname{orb}(y;f)\neq\emptyset,\text{ where }x,y\in X.$ (4)
@@
Proof.
The proofs of being reflexive and symmetric are immediate.
To prove transitivity, let $x,y$ and $z$ be in $X$ satisfying $x\sim y$ and $y\sim z$. Then there are non-negative integers $m,n,m^{\prime}$ and $n^{\prime}$ such that
$f^{m}(x)=f^{n}(y)\text{ and }f^{m^{\prime}}(y)=f^{n^{\prime}}(z).$ (5)
If $n\geq m^{\prime}$ then
$f^{m}(x)$ $=f^{n}(y)$
$=f^{n-m^{\prime}}(f^{m^{\prime}}(y))$
$=f^{n-m^{\prime}}(f^{n^{\prime}}(z))$
$=f^{n-m^{\prime}+n^{\prime}}(z).$
Thus, $\operatorname{orb}(x;f)\cap\operatorname{orb}(z;f)\neq\emptyset$, that is, $x\sim z$. Similarly, it holds that $x\sim z$ if $n<m^{\prime}$. Hence, $\sim$ is an equivalence relation. ∎
@@ 3 Hilbert spaces and operator theory
$\mathbb{K}$ will denote either the real field, $\mathbb{R}$, or the complex field, $\mathbb{C}$. For any set $X$, the Kronecker delta on $X$, denoted by $\delta$, is defined as follow: For $x,y\in X$,
$\delta_{x,y}=\begin{cases}1,&x=y,\\ 0,&x\neq y.\end{cases}$ (6)
@@ 3.1 Hilbert spaces
$\operatorname{d}(x,y)=\left\lvert\left\langle x-y,x-y\right\rangle\right\rvert^{1/2},\text{ where }x,y\in\mathcal{H}.$ (7)
@@
=$\mathcal{M}^{\bot}=\{y\in\mathcal{H}\mid\left\langle x,y\right\rangle=0,\forall x\in\mathcal{M}\}.$ (8)
Theorem 3.1.3 ([4, Chapter I, Theorem 2.5, Theorem 2.6], Orthogonal decomposition).
@@
Definition 3.1.4 (Completely orthogonal normal systems).
Let $\{e_{\lambda}\}_{\lambda\in\Lambda}\subseteq\mathcal{H}$ be a set, where $\Lambda$ is an index set. The net $\{e_{\lambda}\}_{\lambda\in\Lambda}$ is a completely orthogonal normal system (abbreviated as C.O.N.S.) for $\mathcal{H}$, when the following conditions hold:
@@
Theorem 3.1.5 ([4, Chapter I, Theorem 4.13, Proposition 4.14], Fourier series).
@@
    If $\{e_{\lambda}\}_{\lambda\in\Lambda}$ is a C.O.N.S. for $\mathcal{H}$, then every $x\in\mathcal{H}$ can be expanded by $\{\langle x,e_{\lambda}\rangle e_{\lambda}\}_{\lambda\in\Lambda}$, i.e., $x$ is a unique accumulation point of $\{\sum_{\lambda\in F}\left\langle x,e_{\lambda}\right\rangle e_{\lambda}\mid F\text{ is a finite subset of }\Lambda\}$.
@@ 3.2 Bounded linear operators
$\left\lVert x\right\rVert=\left\lvert\left\langle x,x\right\rangle\right\rvert^{1/2},\text{ where }x\in\mathcal{H},$ (9)
@@
$\sup\{\left\lVert Tx\right\rVert\mid x\in\mathcal{H},\left\lVert x\right\rVert\leq 1\}<+\infty.$ (10)
@@
For example, identity map on $\mathcal{H}$ is a bounded linear operator. We will denote this map by $I$.
@@
$\left\langle Tx,y\right\rangle=\left\langle x,T^{*}y\right\rangle$ (11)
@@
For every $T\in\mathcal{B}(\mathcal{H})$, the adjoint of $T$ exists uniquely ([4, Chapter I, Theorem 2.2]). It is easy to see that $I^{*}=I$. More generally, there are many operators $T\in\mathcal{B}(\mathcal{H})$ satisfying that $T^{*}=T$.
@@
$\left\lVert Tx\right\rVert=\left\lVert x\right\rVert$ (12)
@@
$\left\lVert Tx\right\rVert=\left\lVert x\right\rVert$ (13)
@@ 3.3 Reducing subspaces
In this subsection, we recall the notion of invariant subspaces and reducing subspaces. Let $T\in\mathcal{B}(\mathcal{H})$ and $\mathcal{M}\subseteq\mathcal{H}$ be a subspace.
@@ 3.4 $C^{*}$-algebras
$\left\lVert T\right\rVert=\sup\{\left\lVert Tx\right\rVert\mid x\in H,\left\lVert x\right\rVert\leq 1\},\text{ where }T\in\mathcal{B}(\mathcal{H}).$ (15)
@@
$(S+T)x=Sx+Tx,\text{ where }x\in\mathcal{H},$ (16)
$(ST)x=S(Tx),\text{ where }x\in\mathcal{H}.$ (17)
@@
$(\alpha T)x=\alpha(Tx),\text{ where }x\in\mathcal{H}.$ (18)
@@
$\left\{\sum_{i=1}^{n}\alpha_{i}T_{i}\mid n\in\mathbb{N},\alpha_{i}\in\mathbb{K},T_{i}\text{ is an element or a product of elements in }\mathcal{A}\right\}.$ (19)
@@
$\mathcal{A}^{\prime}=\{S\in\mathcal{B}(\mathcal{H})\mid TS=ST(\forall T\in\mathcal{A})\}.$ (20)
@@
$\overline{\mathcal{A}x}=\overline{\{Tx\mid T\in\mathcal{A}\}}=\mathcal{H}.$ (21)
Proposition 3.4.5 ([3, Chapter 2, Proposition 14.3], Sepatrating points).
@@
    $x$ is a separating point for $\mathcal{A}^{\prime}$, i.e., if $T\in\mathcal{A}^{\prime}$ and $Tx=0$, then $T=0$.
@@
Theorem 3.4.6 ([3, Chapter 5, Theorem 32.6], Irreducibility).
@@ 4.1 Formulation by a single operator
$Te_{n}=e_{f(n)},\text{ where }n\in\mathbb{N},$ (22)
@@
$T\left(\sum_{n=1}^{N}x_{n}e_{n}\right)=\sum_{n=1}^{N}x_{n}Te_{n},\text{ where }N\in\mathbb{N},\text{ }x_{n}\in\mathbb{C}.$ (23)
For $x=\sum_{n\in\mathbb{N}}x_{n}e_{n}\in\mathcal{H}$, we can compute the norm of $x$ as $\left\lVert x\right\rVert=\left(\sum_{n\in\mathbb{N}}\left\lvert x_{n}\right\rvert^{2}\right)^{1/2}$. Notice that $f^{-1}(n)$ includes at most two elements for any $n\in\mathbb{N}$ and so $\left\lvert\sum_{m\in f^{-1}(n)}x_{m}\right\rvert^{2}\leq(\sum_{m\in f^{-1}(n)}\left\lvert x_{m}\right\rvert)^{2}\leq 2\sum_{m\in f^{-1}(n)}\left\lvert x_{m}\right\rvert^{2}$. Since
@@
$Tx=\sum_{n\in\mathbb{N}}x_{n}Te_{n}=\sum_{n\in\mathbb{N}}x_{n}e_{f(n)}=\sum_{n\in\mathbb{N}}\left(\sum_{m\in f^{-1}(n)}x_{m}\right)e_{n},$ (24)
@@
$\overline{C^{*}(T)e_{n}}\subseteq\overline{\operatorname{span}\{e_{m}\mid m\in\mathbb{N},m\sim n\}}.$ (25)
@@
When $e_{n}$ is a cyclic vector for $C^{*}(T)$, $\mathcal{H}=\overline{C^{*}(T)e_{n}}\subseteq\overline{\operatorname{span}\{e_{m}\mid m\in\mathbb{N},m\sim n\}}$ by Lemma 4.1.1. It follows that $m\sim n$ for every $m\in\mathbb{N}$. Therefore, $1\sim n$ and then, $1\sim m$ for every $m\in\mathbb{N}$. Since $1$ is a periodic point of $f$, $1\in\operatorname{orb}(m;f)$. Thus, we can obtain the conclusion. ∎
@@ 4.2 Formulation by two operators
$T_{1}e_{n}$ =$=\begin{cases}e_{3n+1},&n:\text{odd},\\ 0,&n:\text{even},\end{cases}$ (26)
$T_{2}e_{n}$ =$=\begin{cases}0,&n:\text{odd},\\ e_{n/2},&n:\text{even}.\end{cases}$ (27)
@@
$\overline{C^{*}(T_{1},T_{2})e_{n}}=\overline{\operatorname{span}\{e_{m}\mid m\in\mathbb{N},m\sim n\}}.$ (28)
@@
Since $T_{1}$ and $T_{2}$ are partial isometries, it holds that $\left\lVert(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}\right\rVert\leq 1$. Thus,
@@
$e_{m}$ $=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}e_{f^{k}(m)}=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}e_{f^{l}(n)}=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}(T_{j_{1}}T_{j_{2}}\cdots T_{j_{l}})e_{n}.$
@@
For $n\in\mathbb{N}$, $T_{1}^{*}T_{1}$ and $T_{2}^{*}T_{2}$ satisfy the following conditions:
$T_{1}^{*}T_{1}e_{n}$ =$=\begin{cases}e_{n},&n:\text{odd},\\ 0,&n:\text{even},\end{cases}$ (31)
$T_{2}^{*}T_{2}e_{n}$ =$=\begin{cases}0,&n:\text{odd},\\ e_{n},&n:\text{even}.\end{cases}$ (32)
@@
00 0 $\neq\left\langle Qe_{1},e_{n}\right\rangle=\left\langle QT_{1}^{*}T_{1}e_{1},e_{n}\right\rangle=\left\langle T_{1}^{*}QT_{1}e_{1},e_{n}\right\rangle=\left\langle QT_{1}e_{1},T_{1}e_{n}\right\rangle$
$=\left\langle Qe_{4},T_{1}e_{n}\right\rangle=\left\langle QT_{2}^{*}T_{2}e_{4},T_{1}e_{n}\right\rangle=\left\langle T_{2}^{*}QT_{2}e_{4},T_{1}e_{n}\right\rangle=\left\langle QT_{2}e_{4},T_{2}T_{1}e_{n}\right\rangle$
$=\left\langle Qe_{2},T_{2}T_{1}e_{n}\right\rangle=\left\langle QT_{2}^{*}T_{2}e_{2},T_{2}T_{1}e_{n}\right\rangle=\left\langle T_{2}^{*}QT_{2}e_{2},T_{2}T_{1}e_{n}\right\rangle=\left\langle PT_{2}e_{2},T_{2}^{2}T_{1}e_{n}\right\rangle$
$=\left\langle Qe_{1},T_{2}^{2}T_{1}e_{n}\right\rangle.$
@@
$Qe_{1}$ $=\sum_{n\in\mathbb{N}}\left\langle Qe_{1},e_{n}\right\rangle e_{n}=\left\langle Qe_{1},e_{1}\right\rangle e_{1}$
and
$\left\lVert Qe_{1}\right\rVert^{2}$ $=\left\lvert\left\langle Qe_{1},e_{1}\right\rangle\right\rvert^{2}=\left\lvert\left\langle Qe_{1},Qe_{1}\right\rangle\right\rvert^{2}=(\left\lVert Qe_{1}\right\rVert^{2})^{2}=\left\lVert Qe_{1}\right\rVert^{4}.$
Since $Qe_{1}\neq 0$, it holds that $\left\lVert Qe_{1}\right\rVert=1$. Then, $Qe_{1}=\left\langle Qe_{1},e_{1}\right\rangle e_{1}=\left\langle Qe_{1},Qe_{1}\right\rangle e_{1}=\left\lVert Qe_{1}\right\rVert^{2}e_{1}=e_{1}$. ∎
@@
For every $C^{*}$-algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$, $\mathcal{A}^{\prime}$ is a von Neumann algebra, which is a special kind of $C^{*}$-algebras. Then $\mathcal{A}^{\prime}$ is the norm closed linear span of its projection (see, e.g. [3, Chapter 2, Proposition 13.3]). Thus,

### v2

@@ Abstract
The Collatz map (or the $13n{+}$ 1-map) $f$ is defined on positive integers by setting $f(n)$ equal to $13n+$ 1 when $n$ is odd and $n/2$ when $n$ is even. The Collatz conjecture states that starting from any positive integer $n$, some iterate of $f$ takes value $1$. In this study, we discuss formulations of the Collatz conjecture by $C^{*}$-algebras in the following three ways: (1) single operator, (2) two operators, and (3) Cuntz algebra. For the $C^{*}$-algebra generated by each of these, we consider the condition that it has no non-trivial reducing subspaces. For (1), we prove that the condition implies the Collatz conjecture. In the cases (2) and (3), we prove that the condition is equivalent to the Collatz conjecture. For similar maps, we introduce equivalence relations by them and generalize connections between the Collatz conjecture and irreducibility of associated $C^{*}$-algebras.
@@ 1 Introduction
Throughout this paper, $\mathbb{N}$ denotes the set of all positive integers and $\left\lvert x\right\rvert$ denotes the absolute value of a number $x$. For a topological space $X$ and its subspace $Y$, $\overline{Y}$ denotes the closure of $Y$ in $X$. For $k\in\mathbb{N}_{>1}$, $\{1,2,\cdots,k\}^{*}$ denotes the disjoint union over $n\in\mathbb{N}$ of $\{1,2,\cdots,k\}^{n}$.
@@ 2.1 Dynamical systems
$P(x)=f^{\tau(x)}(x),\text{ where }x\in\Sigma^{\prime}.$
@@ 2.2 The Collatz conjecture
$f(n)=\begin{cases}3n+1,&n:\text{odd},\\ n/2,&n:\text{even}.\end{cases}$
@@ 2.3 Equivalence relations
We define a binary relation on $X$ which is called the orbit equivalence relation by $f$ as follows: For $x,y\in X$,
$x\sim y\iff\operatorname{orb}(x;f)\cap\operatorname{orb}(y;f)\neq\emptyset,$
in other words, there exist $n,m\in\{0\}\cup\mathbb{N}$ such that $f^{n+k}(x)=f^{m+k}(y)$ for all $k\in\{0\}\cup\mathbb{N}$.
@@ 3 Hilbert spaces and operator theory
$\mathbb{K}$ will denote either the real field, $\mathbb{R}$, or the complex field, $\mathbb{C}$. For any set $X$, the Kronecker delta on $X$, denoted by $\delta$, is defined as follows: For $x,y\in X$,
$\delta_{x,y}=\begin{cases}1,&x=y,\\ 0,&x\neq y.\end{cases}$
@@ 3.1 Hilbert spaces
$\operatorname{d}(x,y)=\left\lvert\left\langle x-y,x-y\right\rangle\right\rvert^{1/2},\text{ where }x,y\in\mathcal{H}.$
@@
=$\mathcal{M}^{\bot}=\{y\in\mathcal{H}\mid\left\langle x,y\right\rangle=0,\forall x\in\mathcal{M}\}.$
Theorem 3.1.3 ([3, Chapter I, Theorem 2.5, Theorem 2.6], Orthogonal decomposition).
@@
Definition 3.1.4 (Completely orthonormal systems).
Let $\{e_{\lambda}\}_{\lambda\in\Lambda}\subseteq\mathcal{H}$ be a set, where $\Lambda$ is an index set. The set $\{e_{\lambda}\}_{\lambda\in\Lambda}$ is a completely orthonormal system (abbreviated as C.O.N.S.) for $\mathcal{H}$, when the following conditions hold:
@@
Theorem 3.1.5 ([3, Chapter I, Theorem 4.13, Proposition 4.14], Fourier series).
@@
    If $\{e_{\lambda}\}_{\lambda\in\Lambda}$ is a C.O.N.S. for $\mathcal{H}$, then every $x\in\mathcal{H}$ can be expanded
    by $\{\langle x,e_{\lambda}\rangle e_{\lambda}\}_{\lambda\in\Lambda}$, i.e., $x$ is a unique accumulation point of the set
    $\{\sum_{\lambda\in F}\left\langle x,e_{\lambda}\right\rangle e_{\lambda}\mid F\text{ is a finite subset of }\Lambda\}$.
@@ 3.2 Bounded linear operators
$\left\lVert x\right\rVert=\left\lvert\left\langle x,x\right\rangle\right\rvert^{1/2},\text{ where }x\in\mathcal{H},$
@@
$\sup\{\left\lVert Tx\right\rVert\mid x\in\mathcal{H},\left\lVert x\right\rVert\leq 1\}<+\infty.$
@@
For example, the identity map on $\mathcal{H}$ is a bounded linear operator. We will denote this map by $I$.
@@
$\left\langle Tx,y\right\rangle=\left\langle x,T^{*}y\right\rangle$
@@
For every $T\in\mathcal{B}(\mathcal{H})$, the adjoint of $T$ exists uniquely ([3, Chapter I, Theorem 2.2]). It is easy to see that $I^{*}=I$. More generally, there are many operators $T\in\mathcal{B}(\mathcal{H})$ satisfying that $T^{*}=T$.
@@
$\left\lVert Tx\right\rVert=\left\lVert x\right\rVert$
@@
$\left\lVert Tx\right\rVert=\left\lVert x\right\rVert$
@@ 3.3 Reducing subspaces
In this subsection, we recall the notion of invariant subspaces and reducing subspaces. Let $T\in\mathcal{B}(\mathcal{H})$ and $\mathcal{M}$ is a subspace of $\mathcal{H}$.
@@ 3.4 $C^{*}$-algebras
$\left\lVert T\right\rVert=\sup\{\left\lVert Tx\right\rVert\mid x\in H,\left\lVert x\right\rVert\leq 1\},\text{ where }T\in\mathcal{B}(\mathcal{H}).$
@@
$(S+T)x=Sx+Tx,\text{ where }x\in\mathcal{H},$
$(ST)x=S(Tx),\text{ where }x\in\mathcal{H}.$
@@
$(\alpha T)x=\alpha(Tx),\text{ where }x\in\mathcal{H}.$
@@
$\left\{\sum_{i=1}^{n}\alpha_{i}T_{i}\mid n\in\mathbb{N},\alpha_{i}\in\mathbb{K},T_{i}\text{ is an element or a product of elements in }\mathcal{A}\right\}.$
@@
$\mathcal{A}^{\prime}=\{S\in\mathcal{B}(\mathcal{H})\mid TS=ST(\forall T\in\mathcal{A})\}.$
@@
$\overline{\mathcal{A}x}=\overline{\{Tx\mid T\in\mathcal{A}\}}=\mathcal{H}.$
Proposition 3.4.5 ([4, Chapter 2, Proposition 14.3], Separating vector).
@@
    $x$ is a separating vector for $\mathcal{A}^{\prime}$, i.e., if $T\in\mathcal{A}^{\prime}$ and $Tx=0$, then $T=0$.
@@
Theorem 3.4.6 ([4, Chapter 5, Theorem 32.6], Irreducibility).
@@ 4.1 Formulation by a single operator
$Te_{n}=e_{f(n)},\text{ where }n\in\mathbb{N},$
@@
$T\left(\sum_{n=1}^{N}x_{n}e_{n}\right)=\sum_{n=1}^{N}x_{n}Te_{n},\text{ where }N\in\mathbb{N},\text{ }x_{n}\in\mathbb{C}.$
For $x=\sum_{n\in\mathbb{N}}x_{n}e_{n}$, we can compute the norm of $x$ as $\left\lVert x\right\rVert=\sqrt{\sum_{n\in\mathbb{N}}\left\lvert x_{n}\right\rvert^{2}}$. Notice that $f^{-1}(n)$ includes at most two elements for any $n\in\mathbb{N}$ and so $\left\lvert\sum_{m\in f^{-1}(n)}x_{m}\right\rvert^{2}\leq(\sum_{m\in f^{-1}(n)}\left\lvert x_{m}\right\rvert)^{2}\leq 2\sum_{m\in f^{-1}(n)}\left\lvert x_{m}\right\rvert^{2}$. Since
@@
$Tx=\sum_{n\in\mathbb{N}}x_{n}Te_{n}=\sum_{n\in\mathbb{N}}x_{n}e_{f(n)}=\sum_{n\in\mathbb{N}}\left(\sum_{m\in f^{-1}(n)}x_{m}\right)e_{n},$
@@
$\overline{C^{*}(T)e_{n}}\subseteq\overline{\operatorname{span}\{e_{m}\mid m\in\mathbb{N},m\sim n\}}.$
@@
When $e_{n}$ is a cyclic vector for $C^{*}(T)$,
$\mathcal{H}=\overline{C^{*}(T)e_{n}}\subseteq\overline{\operatorname{span}\{e_{m}\mid m\in\mathbb{N},m\sim n\}}$
by Lemma 4.1.1. It follows that $m\sim n$ for every $m\in\mathbb{N}$. Therefore, $1\sim n$ and then, $1\sim m$ for every $m\in\mathbb{N}$. Since $1$ is a periodic point of $f$, $1\in\operatorname{orb}(m;f)$. Thus, we can obtain the conclusion. ∎
@@ 4.2 Formulation by two operators
$T_{1}e_{n}$ =$=\begin{cases}e_{3n+1},&n:\text{odd},\\ 0,&n:\text{even},\end{cases}$
$T_{2}e_{n}$ =$=\begin{cases}0,&n:\text{odd},\\ e_{n/2},&n:\text{even}.\end{cases}$
@@
$\overline{C^{*}(T_{1},T_{2})e_{n}}=\overline{\operatorname{span}\{e_{m}\mid m\in\mathbb{N},m\sim n\}}.$
@@
Since $T_{1}$ and $T_{2}$ are partial isometries, $\left\lVert(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}\right\rVert\leq 1$. Thus,
@@
$e_{m}$ $=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}e_{f^{k}(m)}=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}e_{f^{l}(n)}$
$=(T_{i_{1}}T_{i_{2}}\cdots T_{i_{k}})^{*}(T_{j_{1}}T_{j_{2}}\cdots T_{j_{l}})e_{n}.$
@@
We remark that $T_{2}^{2}T_{1}e_{1}=e_{1}$ and $e_{1}=T_{1}^{*}(T_{2}^{*})^{2}$.
@@
$0\neq\left\langle Qe_{1},e_{n}\right\rangle=\left\langle QT_{1}^{*}(T_{2}^{*})^{2}e_{1},e_{n}\right\rangle=\left\langle T_{1}^{*}(T_{2}^{*})^{2}Qe_{1},e_{n}\right\rangle=\left\langle Qe_{1},T_{2}^{2}T_{1}e_{n}\right\rangle.$
@@
$Qe_{1}$ $=\sum_{n\in\mathbb{N}}\left\langle Qe_{1},e_{n}\right\rangle e_{n}=\left\langle Qe_{1},e_{1}\right\rangle e_{1}=\left\lVert Qe_{1}\right\rVert^{2}e_{1}.$
Since $Qe_{1}\neq 0$, it holds that $\left\lVert Qe_{1}\right\rVert=1$ and $Qe_{1}=e_{1}$. ∎
@@
For every $C^{*}$-algebra $\mathcal{A}\subseteq\mathcal{B}(\mathcal{H})$, $\mathcal{A}^{\prime}$ is a von Neumann algebra, which is a special kind of $C^{*}$-algebras. Then $\mathcal{A}^{\prime}$ is the norm closed linear span of its projection (see, e.g. [4, Chapter 2, Proposition 13.3]). Thus,